# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

### v1

@@ Appendix B Comparison of Syntax Elements in Python, C++, and Java
Table 7: Comparison of five types of syntax elements across Python, C++, and Java.
Appendix C Trade-off Between Correct and Detect
We analyze the trade-off between pass@1 and AUROC based on watermarking strength by adjusting two parameters: $\gamma$, which controls the ratio between the green list and red list, and $\delta$, which determines how often the model selects tokens from the green list. Figure 3 shows the results in a bubble chart, revealing a clear trade-off—improving one metric often reduces the other.
The bottom-right section of each subfigure highlights larger light-blue bubbles, where high delta and low gamma values produce higher AUROC but lower pass@1. A smaller gamma shrinks the green list, limiting the token options available during generation. At the same time, a higher delta increases the likelihood of selecting tokens from the green list. This combination strengthens the watermark signal and raises AUROC but generates less natural code, reducing pass@1.
Achieving the right balance between code quality and watermark detectability requires careful tuning of gamma and delta. Maintaining code quality and correctness remains crucial when embedding watermarks, so relying on both a low gamma and a high delta together creates undesirable results. The areas shaded in green highlight points where pass@1 and AUROC achieve a proper balance. The double circles outlined in black represent the performance reported in Table 2. For the MBPP+ dataset, these points correspond to $\delta$ = 1.0 and $\gamma$ = 0.5, while for the HumanEval+ dataset, they correspond to $\delta$ = 0.5 and $\gamma$ = 0.5. The trends in the green-shaded region suggest that using a lower delta value generally results in a better balance between pass@1 and AUROC performance.
Figure 3: The tradeoff between Correct and Detect. We analyze the impact of the green list ratio $\gamma$ (indicated by marker color) and the logit adjustment factor $\delta$ (indicated by marker size). Each point represents the functional correctness of code watermarked using the STONE (Y-axis) and the performance of classifying it against human-written code (X-axis). The green-shaded area highlights points where Correct and Detect are well balanced. Double circles with an outer black ring mark the performance points reported in Table 2.

### v2

@@ Appendix B Performance Analysis of CodeIP
Table 7: Comparison of five categories of syntax elements across Python, C++, and Java.
Appendix C Experimental Results on Llama-3.1-8B
Tables 8 and 9 summarize the experimental results obtained with the Llama-3.1-8B model on MBPP+ and HumanEval+. STONE achieves the highest STEM scores across all datasets and weight settings, demonstrating that the non-syntax watermarking strategy generalizes well to larger models. EWD shows slightly higher detectability on HumanEval+, which appears to result from an interaction between its uniform insertion strategy and the sharper probability distributions of Llama-3.1. In this setting, the cumulative bias introduced by EWD is more effectively reinforced across long and diverse sequences, leading to a small gain in detectability. Such an effect is not observed under Qwen2.5-Coder-7B, whose flatter distributions yield comparable results between EWD and SWEET. Overall, STONE preserves stronger correctness and achieves a more balanced trade-off, resulting in the highest overall STEM scores.
Across different weight configurations, STONE maintains stable performance even when correctness or imperceptibility is prioritized. The method remains competitive under detectability-oriented settings without functional degradation. This consistency confirms that syntax-aware watermarking effectively preserves code quality while sustaining robust watermark signals.
Method MBPP+ HumanEval+ Correct. Detect. Imperc. Correct. Detect. Imperc. KGW 0.414 0.831 0.953 0.323 0.507 0.989 EWD 0.414 0.931 0.953 0.323 0.755 0.989 SWEET 0.404 0.927 0.968 0.354 0.711 0.978 STONE 0.446 0.945 0.954 0.372 0.741 0.983
Table 8: Experimental results on MBPP+ and HumanEval+ (Llama-3.1-8B): Correctness, Detectability, and Imperceptibility.
  Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ KGW EWD SWEET STONE KGW EWD SWEET STONE $(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.733 0.766 0.766 0.782 0.606 0.689 0.681 0.699 (0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.653 0.678 0.676 0.698 0.535 0.598 0.599 0.617 (0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.757 0.807 0.806 0.823 0.582 0.706 0.688 0.709 (0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.788 0.813 0.817 0.825 0.702 0.764 0.755 0.770
Table 9: STEM scores on MBPP+ and HumanEval+ (Llama-3.1-8B) under different weight settings $(\alpha,\beta,\zeta)$. The left column lists weights for correctness, detectability, and imperceptibility. Bold indicates the highest score for each weight configuration.
Entropy Threshold MBPP+ HumanEval+ HEP-C++ HEP-Java
Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%)
0.7 26.03 12.82 35.28 12.84 30.10 12.99 26.60 11.98
0.8 21.40 12.59 32.67 12.24 25.93 12.35 22.47 11.06
0.9 19.64 11.39 28.98 12.60 22.65 11.79 18.52 11.68
1.0 16.59 11.39 24.84 11.83 18.96 10.56 16.13 11.15
1.1 14.94 10.30 23.22 11.81 15.71 9.82 14.39 10.91
Table 10: Token selection statistics of SWEET across four datasets. Selected Tokens (%) denotes the proportion of tokens exceeding the entropy threshold out of all tokens, and Syntax Tokens (%) represents the percentage of syntax-related tokens among those selected. Bold values indicate the optimal entropy threshold used in SWEET.
  Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ HEP-C++ HEP-Java KGW EWD SWEET STONE KGW EWD SWEET STONE KGW EWD SWEET STONE KGW EWD SWEET STONE $(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.775 0.819 0.787 0.848 0.694 0.763 0.754 0.781 0.730 0.750 0.735 0.780 0.642 0.675 0.631 0.715 (0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.706 0.739 0.716 0.778 0.664 0.716 0.709 0.732 0.692 0.706 0.697 0.741 0.578 0.603 0.577 0.648 (0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.789 0.856 0.807 0.881 0.651 0.755 0.743 0.780 0.703 0.733 0.711 0.768 0.618 0.668 0.618 0.716 (0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.830 0.863 0.838 0.883 0.767 0.810 0.808 0.830 0.796 0.811 0.796 0.833 0.730 0.755 0.699 0.781
Table 11: STEM scores under different weight settings $(\alpha,\beta,\zeta)$. The left column lists weights for correctness, detectability, and imperceptibility. Bold indicates the best score in each row.
Appendix D Comprehensive Evaluation across STEM Weight Configurations
Table 11 reports representative STEM variants using equal weights and three dimension-focused settings. We further explore all valid weight configurations where $\alpha,\beta,\zeta\in\{0.0,0.1,\dots,1.0\}$ and $\alpha+\beta+\zeta=1$, resulting in $66$ combinations. For each configuration, we compute the weighted STEM score and record the method with the highest value. Across all datasets, STONE performs consistently well throughout the entire weight space, ranking first in $97.0\%$ of configurations on MBPP+, $90.9\%$ on HumanEval+, $98.5\%$ on HumanEvalPack-C++, and $95.5\%$ on HumanEvalPack-Java. This consistency indicates that the advantage of STONE does not depend on specific weight choices but generalizes across diverse evaluation preferences, demonstrating robust performance under varying priorities of correctness, detectability, and imperceptibility.
Appendix E Syntax Token Coverage in SWEET Watermarking
Table 10 presents an analysis of token selection by SWEET using two measures: (1) the proportion of selected tokens exceeding the entropy threshold among all generated tokens, and (2) the proportion of syntax-related tokens among those selected. We use the optimal settings ($\gamma$ = 0.25, $\delta$ = 3.0, entropy threshold = 0.9)) and evaluate threshold values in the range [0.7, 0.8, 0.9, 1.0, 1.1]. The results show an inverse relationship between the threshold and token coverage: lower thresholds select more tokens overall but include more syntax elements, whereas higher thresholds reduce syntax token inclusion. This pattern shows that SWEET effectively targets high-entropy regions but does not completely avoid syntax-related tokens.
Even under the optimal configuration, a notable portion of syntax tokens remains among those selected for watermarking. In the HumanEval+ dataset, syntax tokens account for 12.6% of all selected tokens, including delimiters (49.29%), whitespace (38.17%), keywords (9.44%), types (3.17%), and operators (2.78%). This indicates that delimiters and whitespace dominate syntax-related selections, which directly affects code execution.
Category Examples
[…]
Types int, set, str
Delimiters . , : , {, }, [, ], (, )
Operators *, %, =, -, **, &, ^, /, +, |
Table 12: Examples of syntax tokens selected for watermarking in SWEET.
Table 12 lists representative examples. Delimiters such as colons (‘:’) and commas (‘,’), essential for code structure, cause parsing errors when modified (e.g., omitting a colon in ‘for i in range(10):’ leads to a failure). Although keywords and operators appear less frequently, their modification can still alter program behavior, for example, replacing ‘True’ with ‘False’ or ‘+’ with ‘-’. Minimizing the inclusion of such syntax-critical tokens can make syntax-aware watermarking more reliable and less disruptive to functionality.
Appendix F Comparison with Non-watermarked Code
Method MBPP+ HumanEval+
No watermark 0.571 0.595
KGW 0.499 0.573
EWD 0.499 0.573
SWEET 0.502 0.574
STONE 0.571 0.587
Table 13: Correctness comparison between non-watermarked and watermarked code. Correctness is measured as pass@1 on MBPP+ and HumanEval+.
We evaluate the functional impact of watermarking by comparing the correctness of code generated with and without watermark insertion. Non-watermarked code defines the upper bound of functional performance. STONE achieves scores nearly identical to this bound, showing that watermark insertion has minimal influence on code execution. Among all watermarking methods, STONE reduces correctness the least, which indicates that it embeds provenance information effectively while preserving the original program behavior.